OpenStudy (anonymous):

The number of endangered species in a jungle decreased from 75 to 30 in one year. What is the percent of decrease in the number of endangered species?

OpenStudy (aihberkhan):

Use this: percent decrease = (old - new) / old, then multiply that by 100 to get your percent

OpenStudy (anonymous):

So what would the new and the old numbers be would it be 30 as the old one and the new one would be 75

OpenStudy (anonymous):

it is 40%

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

Liam u need to explain

OpenStudy (anonymous):

ok

OpenStudy (study312):

75 - 30 = 45. 45 ÷ 75 = 0.6, 0.6 x 100 = 60. I think it's 60%

OpenStudy (anonymous):

well 3 is 1/25 of 75 thus making 30 40% of 75 or you can do 75 x 0.40 which = 30

OpenStudy (anonymous):

oh nvm

OpenStudy (anonymous):

dang it I was wrong since you decreasing 45

OpenStudy (anonymous):

So I DId 30/70 and got .40 then that times 100 equals 40

OpenStudy (anonymous):

I mean 30/75

OpenStudy (study312):

that can't be right, it's 75 - 30

OpenStudy (study312):

which equals 45

OpenStudy (anonymous):

I can answer it if it was 45 which would be 60%

OpenStudy (anonymous):

Lemme call for back up

OpenStudy (anonymous):

which one is it subtracting 45 or 30?

OpenStudy (study312):

No it's 60%

OpenStudy (anonymous):

yea 60% look up

OpenStudy (anonymous):

I just said that if it wasn't 30

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

cause I realized that it could be something else

OpenStudy (study312):

Liam it's 75 - 30 = 45 45 ÷ 75 = 0.6 0.6 x 100 = 60

OpenStudy (anonymous):

yes dude

OpenStudy (anonymous):

look up

OpenStudy (anonymous):

I said that already

OpenStudy (study312):

Liam you REALLY think it is 40?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I made a mistake

OpenStudy (anonymous):

now look up and you will see I said it was 60% after realizing what mistake I made

OpenStudy (anonymous):

go up 17 replies and you will see what I said

OpenStudy (anonymous):

Thanks guys for helping it's ok if you made a mistake you both helped as a team thanks for the effort and work I Understand now........................

OpenStudy (anonymous):

np

OpenStudy (anonymous):

anything else you got cause I am trying to hit 60 smart score today

OpenStudy (anonymous):

Oh ok yass lemme get tem................